Find all possible Taylor's series and Laurent series expansions of f(z) = (2z-3) /(z-2) (z-1) about z=0?

1
Expert's answer
2022-02-11T11:14:56-0500

Solution:

Consider "f(z)=\frac{z-1}{z^{2}-2 z-3}=\frac{z-1}{(z-3)(z+1)}"

Using partial fraction we get

 "\frac{z-1}{(z-3)(z+1)}=\frac{a}{z-3}+\frac{b}{z+1}"

i.e. z-1=a(z+1)+b(z-3)

put z=3 , we get "a=\frac{1}{2}"

put z=-1 , we get "b=\frac{1}{2}"

"\frac{z-1}{(z-3)(z+1)}=\frac{\frac{1}{2}}{z-3}+\frac{\frac{1}{2}}{z+1}"

Hence f(z) is not analytic at z=-1 and z=3

"\therefore" f(z) is analytic in

"* |z|<1 
\\
* 1<|z|<3 
\\
* |z|>3"

(i)

"|z|<1 

 \\\because|z|<1,|z|<3"

"\begin{aligned}

&\therefore f(z)=\frac{\frac{1}{2}}{z-3}+\frac{\frac{1}{2}}{z+1} \\

&=\frac{1}{2} \cdot \frac{1}{-3} \cdot \frac{1}{1-\frac{z}{3}}+\frac{1}{2} \frac{1}{1+z} \\

&=-\frac{1}{6}\left(1-\frac{z}{3}\right)^{-1}+\frac{1}{2}(1+z)^{-1}

\end{aligned}"

"\begin{aligned}

&=-\frac{1}{6}\left(1+\frac{z}{3}+\left(\frac{z}{3}\right)^{2}+\left(\frac{z}{3}\right)^{3}+\ldots \ldots\right)+\frac{1}{2}\left(1-z+z^{2}-z^{3}+\ldots \ldots\right) \\

&f(z)=\frac{1}{3}-\frac{5}{9} z+\frac{13}{27} z^{2}-\ldots \ldots

\end{aligned}"


Above is required Taylor's series

(ii) 1<|z|<3

We get "\left|\frac{1}{z}\right|<1"  and "\left|\frac{z}{3}\right|<1"

"\begin{aligned}

&\therefore f(z)=\frac{\frac{1}{2}}{z-3}+\frac{\frac{1}{2}}{z+1} \\

&=\frac{1}{2} \cdot \frac{1}{-3} \cdot \frac{1}{1-\frac{z}{3}}+\frac{1}{2} \cdot \frac{1}{z} \cdot \frac{1}{1+\frac{1}{z}} \\

&=\frac{1}{-6} \cdot \frac{1}{1-\frac{z}{3}}+\frac{1}{2 z} \cdot \frac{1}{1+\frac{1}{z}} \\

&=-\frac{1}{6}\left(1-\frac{z}{3}\right)^{-1}+\frac{1}{2 z}\left(1+\frac{1}{z}\right)^{-1} \\

&=-\frac{1}{6}\left(1+\frac{z}{3}+\left(\frac{z}{3}\right)^{2}+\left(\frac{z}{3}\right)^{3}+\ldots \ldots\right)+\frac{1}{2 z}\left(1-\left(\frac{1}{z}\right)+\left(\frac{1}{z}\right)^{2}-\left(\frac{1}{z}\right)^{3}+\ldots\right)^{2}+\frac{1}{2}\left(\frac{1}{z}-\left(\frac{1}{z}\right)^{2}+\left(\frac{1}{z}\right)^{3}-\left(\frac{1}{z}\right)^{4}+\ldots\right. \\

&f(z)=-\frac{1}{6}\left(1+\frac{z}{3}+\left(\frac{z}{3}\right)^{2}+\left(\frac{z}{3}\right)^{3}+\ldots\right)

\end{aligned}"

This is required Laurent's series

(iii) |z|>3

"\begin{aligned}

&\because|z|>3, \text { clearly, }|z|>1 \\

&\therefore\left|\frac{z}{3}\right|>1 \text { and }|z|>1 \\

&\text { i.e. }\left|\frac{3}{z}\right|<1 \text { and }\left|\frac{1}{z}\right|<1 \\

&\therefore f(z)=\frac{\frac{1}{2}}{z-3}+\frac{\frac{1}{2}}{z+1} \\

&=\frac{1}{2 z} \cdot \frac{1}{1-\frac{3}{z}}+\frac{1}{2 z} \cdot \frac{1}{1+\frac{1}{z}} \\

&=\frac{1}{2 z}\left(1-\frac{3}{z}\right)^{-1}+\frac{1}{2 z}\left(1+\frac{1}{z}\right)^{-1} \\

&=\frac{1}{2 z}\left(1+\left(\frac{3}{z}\right)+\left(\frac{3}{z}\right)^{2}+\left(\frac{3}{z}\right)^{3}+\ldots\right)+\frac{1}{2 z}\left(1-\left(\frac{1}{z}\right)+\left(\frac{1}{z}\right)^{2}-\left(\frac{1}{z}\right)^{3}+\ldots .\right) \\

&=\frac{1}{z}+\frac{1}{z^{2}}+\frac{5}{z^{3}}+\frac{13}{z^{4}}+\ldots \cdots \cdots \\

&f(z)=\frac{1}{z}+\frac{1}{z^{2}}+\frac{5}{z^{3}}+\frac{13}{z^{4}}+\ldots \ldots \ldots .

\end{aligned}"

This is the required Laurent's series